OpenStudy (anonymous):

Solve for t. c = rt + k A. t = c -k/r B. t = c+k/r C. t = c/r-k D. t = c-r/k

OpenStudy (anonymous):

first you subtract k to the other side then divide r \[ c = rt + k \rightarrow c-k=rt \] \[\frac{ c-k }{ r }=t\]